A sample of 150 students were chosen at random from all grade 11 students in a private school in Pampanga. The sample indicated that 75% of them were in favor of having an educational trip. Find a) 95% and b) 99% confidence intervals for the proportion of all students who are in favor of having an educational trip. 


1
Expert's answer
2021-05-07T09:05:00-0400

"\text{0.95 confidence interval}\\
L.L=\hat p - z_\frac{\alpha} {2} \sqrt{\frac{\hat p (1-\hat p) }{n}} \\
=0.75 - 1.96\sqrt{\frac{0.75(0.25) }{150}} \approx 0.681\\
U.L=\hat p +z_\frac{\alpha} {2} \sqrt{\frac{\hat p (1-\hat p) }{n}} \\
0.75 +1.96\sqrt{\frac{0.75(0.25) }{150}} \approx 0.819\\
\text{0.99 confidence interval}\\
L.L=\hat p - z_\frac{\alpha} {2} \sqrt{\frac{\hat p (1-\hat p) }{n}} \\
=0.75 - 1.65\sqrt{\frac{0.75(0.25) }{150}} \approx 0.692\\
U.L=\hat p +z_\frac{\alpha} {2} \sqrt{\frac{\hat p (1-\hat p) }{n}} \\
0.75 +1.65\sqrt{\frac{0.75(0.25) }{150}} \approx 0.808\\"